# An Ecosystem for Ontology Interoperability

### v3

@@ Abstract
Ontology interoperability is one of the complicated issues that restricts the use of ontologies in knowledge graphs (KGs). Different ontologies with conflicting and overlapping concepts make it difficult to design, develop, and deploy an interoperable ontology for downstream tasks. We propose an ecosystem for ontology interoperability. The ecosystem employs three state-of-the-art semantic techniques in different phases of the ontology engineering life cycle: ontology design patterns (ODPs) in the design phase, ontology matching and versioning (OM&OV) in the develop phase, and ontology-compliant knowledge graphs (OCKGs) in the deploy phase, to achieve better ontology interoperability and data integration in real-world applications. A case study of sensor observation in the building domain validates the usefulness of the proposed ecosystem.
Index Terms:
ontology, data interoperability
@@ I Introduction
In this paper we review three state-of-the-art semantic web techniques: ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and ontology-compliant knowledge graphs (OCKGs). We compare the differences and connections between these three techniques and explore their roles in ontology engineering to improve interoperability. We propose an ecosystem for ontology interoperability and use a case study of sensor observation in the building domain to showcase how these three techniques can complement each other’s strengths and mitigate their weaknesses within the ontology engineering life cycle.
@@ II-A Ontology Design Patterns
The goal of ontology design patterns (ODPs) is to explore common and shared ontology fragments between different ontologies [2]. Fig. 1 illustrates that ODPs use a star-based matching schema. Once the ODP is determined, each subsequent ontology will conform to this ODP, and the ODP serves as a high-level conceptual model to guide the design of these new ontologies.
@@
However, ODPs also have drawbacks. Firstly, ODPs are conceptual models and do not contain details of the actual class or property. This means that it is not possible to control the terminologies, and ontology creators are still free to use their own choices. Secondly, ODPs are commonly hand-crafted on the basis of existing ontologies. They are static and tend not to capture changes over time. Thirdly, the choice of ODPs is entirely reliant on one’s experience and preferences; the chosen ODP may be overwhelming, overlapping, or scarce for the implemented KG.
@@ II-B Ontology Matching and Versioning
OM&OV are typically resource-intensive. The complexity of pairwise mapping is the product of the number of entities in each of the two ontologies. The pairwise mapping process also needs to be repeated multiple times if multiple ontologies exist. Moreover, OM&OV tasks focus on either conceptual-level matching (i.e. TBox-to-TBox matching) or instance-level matching (i.e. ABox-to-ABox matching), and less attention has been paid to the compliance between conceptual models and data instances (i.e. TBox-to-ABox matching).
II-C Ontology-Compliant Knowledge Graphs
Ontology-compliant knowledge graphs (OCKGs) use a pay-as-you-go model to find the most cost-effective ontology fragment that matches the given KG(s) [5]. Fig. 3 shows that OCKGs comply with a star-based matching schema. Whenever ontologies comply with KG(s), their reference mappings can be derived from different ontology classes or properties pointing to the same KG entities.
Figure 3: Ontology-compliant knowledge graphs (OCKGs).
Building OCKGs is not easy. To determine which candidate ontology fits the KG(s) better, the KG needs to be modelled with each candidate. This process itself is essentially building an alignment of the candidate ontologies, informed by the KG under study.
@@ III An Ecosystem
Fig. 4 illustrates an overview of an ecosystem for ontology interoperability. Similarly to software engineering, ontology engineering involves three main phases: design, develop, and deploy. We employ ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and ontology-compliant knowledge graphs (OCKGs) in different phases of the ontology engineering life cycle to achieve interoperability.
@@ IV-A Design Phase
Ontologies can be built from scratch, but such ontologies are likely to have more customised concepts and limited interoperability. It is commonly recommended to reuse pre-existing ODPs. Fig. 5 illustrates that four classical ODPs can be used in this case to model the example of sensor observation from different perspectives. While all four of these ODPs can effectively capture the main concepts, significant differences in perspective need to be considered, as shown in the following.
  1. (1)
@@
  1. 1.
    The definition of the concepts can be different. For example, temperature is defined as a property of the room in both Observation Pattern and Measurement Pattern, but in the Event Pattern, it is considered a quality measure of the room. Activity Pattern defines the temperature as a member of the room.
  2. 2.
    The levels of expressiveness are different. Unlike Observation Pattern, Measurement Pattern, and Activity Pattern, which can only describe a sensor that is installed in the room, the Event Pattern can distinguish whether it is a structural attachment (i.e. isPartOf) or only a spatial presence (i.e. isLocatedIn).
Figure 5: Modelling using different ODPs.
@@ IV-B Develop Phase
We assume that the Observation Pattern is chosen in the design phase and implemented in two ontologies, SOSA and the original SSN [8, 12] (also called the old SSN in the W3C standard [13]). As illustrated in Fig. 6, we can see that SOSA and SSN are very similar because they are built from the same ODPs. However, they are not exactly the same. The OM task is to find aligned concepts between two different ontologies. SOSA and an updated SSN were co-developed, with a published alignment between them. Sosa was more of a module of the updated SSN (for which there were other modules too). If we consider SOSA to be an updated and reduced version of SSN, then this OM task is also an OV task. We use OM&OV here because OM and OV can be unified with minor modifications [14]. OM&OV tasks can be classified into two main categories.
@@ IV-C Deploy Phase
Now we have developed SOSA and SSN, along with their reference alignment. However, it is hard to determine which ontology to use without applying the real-world KGs. Some concepts used in a KG may not have the corresponding classes and properties in the ontology. Therefore, we apply these two ontologies in the KGs and construct the OCKGs for evaluation. Fig. 7 shows the OCKGs based on SOSA and SSN. We can see that both ontologies are able to capture the general concepts of this sensor observation; however, they fail to capture two specific concepts within the given KG.
  1. 1.
@@
To model these two new concepts and adequately integrate them with the other parts of a KG, we need to extend the ODP with the Stimulus-Sensor-Observation Pattern from SSN and the Sampling Pattern from SOSA. We also make several design choices between aligned mappings (retrieved from [13, 15, 16]).
  1. 1.
@@
Figure 7: Build OCKGs using SOSA and SSN.
Fig. 8 shows the updated ontology and the OCKG built on the updated ontology. Apparently, the updated ontology is more suitable for the provided KG instance. In the real world, an ontology can be deployed in multiple KGs, and the procedure described above can be repeated multiple times. It is also possible to choose other alternative ODPs. For example, the class “BuildingSpace” in the Measurement Pattern is more suitable to describe a room, and the property “hasLocation/isPartOf” in the Event Pattern is more appropriate to describe the link between the sensor and the room.
@@ IV-D Implications
Well-defined ODPs can be used to create reusable ontology fragments, thereby reducing the complexity of OM&OV tasks. As a complement to ODPs, OM&OV can align the concepts that are not captured by ODPs. On the other hand, while OCKGs are useful tools for validating ontology interoperability and evolving ODPs, the construction of OCKGs is heavily based on the precise mapping produced by OM&OV systems. OM&OV is a critical middleware between ODPs and OCKGs.
V Discussion & Conclusion
Ontology interoperability has been a long-standing issue studied in the Semantic Web community. Since the establishment of the International Semantic Web Conference (ISWC), workshops on ontology design and patterns (WOP) and ontology matching (OM) have been the two longest-running workshop proceedings. Ontology-compliant knowledge graphs (OCKGs) aim to explore the interoperability between an ontology and its complementary KGs.
While there is no preference for one technique over another, these three techniques play different roles in an ecosystem for ontology interoperability. We believe that employing ODPs in the design phase, utilising OM&OV in the develop phase, and constructing OCKGs in the deploy phase could significantly improve interoperability throughout the ontology engineering life cycle.

### v5

@@ Abstract
Ontology interoperability is one of the complicated issues that restricts the use of ontologies in knowledge graphs (KGs). Different ontologies with conflicting and overlapping concepts make it difficult to design, develop, and deploy an interoperable ontology for downstream tasks. We propose an ecosystem for ontology interoperability. The ecosystem employs three state-of-the-art semantic techniques in different phases of the ontology engineering life cycle: ontology design patterns (ODPs) in the design phase, ontology matching and versioning (OM&OV) in the develop phase, and data-driven ontology validation (DOVA) in the deploy phase, to achieve better ontology interoperability and data integration in real-world applications. A case study of sensor observation in the building domain validates the usefulness of the proposed ecosystem.
@@ I Introduction
In this paper, we review three state-of-the-art semantic web techniques: ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and ontology-compliant knowledge graphs (OCKGs). We compare the differences and connections between these three techniques and explore their roles in ontology engineering to improve interoperability. We propose an ecosystem for ontology interoperability and use a case study of sensor observation in the building domain to showcase how these three techniques can complement each other’s strengths and mitigate their weaknesses within the ontology engineering life cycle.
@@ II-A Ontology Design Patterns
The goal of ontology design patterns (ODPs) is to identify common ontology fragments across different ontologies [2]. Fig. 1 illustrates that ODPs use a star-based matching schema. Once the ODP is determined, each subsequent ontology will conform to this ODP, and the ODP serves as a high-level conceptual model to guide the design of these new ontologies.
@@
However, ODPs also have drawbacks. Firstly, ODPs are conceptual models and do not contain details of the actual class or property. This means that it is not possible to control the terminologies, and ontology creators are still free to use their own choices. Secondly, ODPs are commonly handcrafted on the basis of existing ontologies. They are static and tend not to capture changes over time. Thirdly, the choice of ODPs is entirely reliant on one’s experience and preferences; the chosen ODP may be overwhelming, overlapping, or scarce for the implemented KG.
@@ II-B Ontology Matching and Versioning
OM&OV are typically resource-intensive. The complexity of pairwise mapping is the product of the number of entities in the two ontologies. The pairwise mapping process also needs to be repeated multiple times if multiple ontologies exist. Moreover, OM&OV tasks focus on either conceptual-level matching (i.e. TBox-to-TBox matching) or instance-level matching (i.e. ABox-to-ABox matching), and less attention has been paid to the compliance between conceptual models and data instances (i.e. TBox-to-ABox matching).
II-C Data-Driven Ontology Validation
Data-driven ontology validation (DOVA) uses KGs to validate the effectiveness of the ontology. Fig. 3 shows that DOVA complies with a star-based matching schema. Whenever ontologies comply with KG(s), their reference mappings can be derived from different ontology classes or properties pointing to the same KG entities.
Figure 3: Data-driven ontology validation (DOVA).
The introduction of ontology-compliant knowledge graphs (OCKGs) is a novel DOVA approach that uses a pay-as-you-go model to find the most cost-effective ontology fragment that matches the given KG(s) [5]. Building OCKGs is not easy. To determine which candidate ontology best fits the KG(s), the KG needs to be modelled for each candidate. This process essentially involves aligning the candidate ontologies, informed by the KG under study.
@@ III An Ecosystem
Fig. 4 illustrates an overview of an ecosystem for ontology interoperability. Similar to software engineering, ontology engineering involves three main phases: design, develop, and deploy. We employ ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and data-driven ontology validation (DOVA) in different phases of the ontology engineering life cycle to achieve interoperability.
@@ IV-A Design Phase
Ontologies can be built from scratch, but such ontologies are likely to have more customised concepts and limited interoperability. It is commonly recommended to reuse pre-existing ODPs. Fig. 5 illustrates that four classical ODPs can be used in this case to model the example of sensor observation from different perspectives. While all four of these ODPs can effectively capture the main concepts, significant differences in perspective need to be considered, as shown below.
  1. (i)
@@
  1. (i)
    The definitions of the concepts can differ. Temperature is defined as a property of the room in both the Observation Pattern and the Measurement Pattern, but in the Event Pattern it is treated as a quality measure of the room. Activity Pattern defines the temperature as a member of the room.
  2. (ii)
    The levels of expressiveness are different. Unlike Observation Pattern, Measurement Pattern, and Activity Pattern, which can only describe a sensor that is installed in the room, the Event Pattern can distinguish between a structural attachment (i.e. isPartOf) and a spatial presence (i.e. isLocatedIn).
@@ IV-B Develop Phase
We assume that the Observation Pattern is chosen in the design phase and implemented in two ontologies, SOSA and the original SSN [8, 12] (also called the old SSN in the W3C standard [13]). As illustrated in Fig. 6, SOSA and SSN are very similar because they are built from the same ODPs. However, they are not exactly the same. The OM task is to identify aligned concepts across two ontologies. SOSA and an updated SSN were co-developed, with a published alignment between them. Sosa was more of a module of the updated SSN (for which there were other modules too). If we consider SOSA to be an updated and reduced version of SSN, then this OM task is also an OV task. We use OM&OV here because OM and OV can be unified with minor modifications [14]. OM&OV tasks can be classified into two main categories.
@@ IV-C Deploy Phase
Now we have developed SOSA and SSN, along with their reference alignment. However, it is hard to determine which ontology to use without applying the real-world KGs. Some concepts used in a KG may not have the corresponding classes and properties in the ontology. Therefore, we apply these two ontologies in the KGs and construct the OCKGs for evaluation. Fig. 7 shows the OCKGs based on SOSA and SSN. We can see that both ontologies can capture the general concepts of this sensor observation; however, they fail to capture two specific concepts in the given KG.
Figure 7: Build OCKGs using SOSA and SSN.
  1. (i)
@@
To model these two new concepts and adequately integrate them with the other parts of a KG, we need to extend the ODP with the Stimulus-Sensor-Observation Pattern from SSN and the Sampling Pattern from SOSA, as well as make several design choices between aligned mappings (retrieved from [13, 15, 16]).
  1. (i)
@@
Fig. 8 shows the updated ontology and the OCKG built on the updated ontology. Apparently, the updated ontology is better suited to the provided KG instance. In the real world, an ontology can be deployed in multiple KGs, and the procedure described above can be repeated multiple times. It is also possible to choose other alternative ODPs. For example, the class “BuildingSpace” in the Measurement Pattern is more suitable to describe a room, and the property “hasLocation/isPartOf” in the Event Pattern is more appropriate to describe the link between the sensor and the room.
@@ IV-D Implications
Well-defined ODPs can be used to create reusable ontology fragments, thereby reducing the complexity of OM&OV tasks. As a complement to ODPs, OM&OV can align concepts that are not captured by ODPs. On the other hand, while OCKGs are useful tools for validating ontology interoperability and evolving ODPs, the construction of OCKGs is heavily based on the precise mapping produced by OM&OV systems. OM&OV is a critical middleware between ODPs and OCKGs.
V Contributions
  1. (i)
    ODPs: We systematically analyse four popular ontologies used in the smart building domain and observe that each building ontology has merits and demerits. We propose building ODPs to capture common and well-agreed concepts and relations in the building domain, including building spaces, equipment and systems, points and measurements [6].
  2. (ii)
    OM&OV: We develop a software system for OM and OV. Agent-OM [17] uses state-of-the-art LLM agents that leverage its novel infrastructure to deliver improved OM performance in a fully-automated and domain-independent manner. Agent-OV [14] evolves from Agent-OM and explicitly targets OV tasks, including new task formulation and evaluation, with a cross-reference mechanism that compiles OM tasks. Agent-OM is used to create an LLM hallucination dataset [18] for the LLM leaderboard and LLM fine-tuning for OM. Agent-OM also participate in the Ontology Alignment Evaluation Initiative (OAEI) [19] 2025 campaign.
  3. (iii)
    DOVA: We introduce OCKGs and define three subtasks under OCKGs: ontology compliance within KG, ontology compliance over KGs, and pattern-based compliance. Examples in the building domain showcase the efficiency of using OCKGs in ontology validation and decision-making [5].
VI Discussion & Conclusion
Ontology interoperability has been a long-standing issue studied in the Semantic Web community. Since the establishment of the International Semantic Web Conference (ISWC), workshops on ontology design and patterns (WOP) and ontology matching (OM) have been the two longest-running workshop proceedings. Data-driven ontology validation aims to explore the interoperability between an ontology and its complementary KGs.
While there is no preference for one technique over another, these three techniques play different roles in an ecosystem for ontology interoperability. We believe that employing ODPs in the design phase, utilising OM&OV in the develop phase, and using DOVA in the deploy phase could significantly improve interoperability throughout the ontology engineering life cycle.